# Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v4

@@ IV Case Studies
We perform a thorough performance analysis of models in Table I on multiple hardware platforms listed in  Table III. The evaluated hardware platforms span workstation and data center-class CPUs and GPUs. We employ PyTorch, ONNX runtime, and TensorRT as deployment flows to investigate their impact on the performance horizon.
Using the aforementioned workload, hardware platforms, and deployment flow, we characterize the GEMM/non-GEMM performance for each model in NonGEMM Bench and investigate the impact of the hardware platform and deployment flow. In addition, we also investigate the impact of widely adopted optimizations (operator fusion and quantization) and show that the non-GEMM performance is an issue even after applying common optimizations.
@@ IV-A Non-GEMM Performance Characterization Results
We first conduct a performance characterization study using PyTorch and present the characterization results on datacenter and workstation platforms (Platforms A and B in Table III) in Figure 5. We focus on those two platforms since some models (e.g, large language models) do not fit into the mobile devices. As expected, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GPU acceleration, on average, from 17.2% to 42.3%. However, we observe each model show different trends, mainly affected by the non-GEMM operator types and the number of GEMM and non-GEMM operators in the original model. We summarize the most time-consuming non-GEMM operators in Table IV from the datacenter class platform (Platform A), which shows the diversity of the dominating non-GEMM operators in each model. We highlight some notable models in each task category and delve into the details of their non-GEMM performance.
@@
IS Task: Maskformer. MaskFormer utilizes Swin Transformer as its backbone, which introduces many memory operators as we discussed for Swin Transformer. As a result, memory operator is the dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 5 (f) and (n).
NLP Task: GPT2. The latencies of non-GEMM operators in GPT2 variants are considerable, which account for 45.0%, on average, as we observe in Figure 5 (h) and (p). The dominant non-GEMM operator is activation, GELU, which required 26.4% of the total latency.
@@
Summary. We observe GPU acceleration significantly increase the percentage of non-GEMM operators and increase the importance of non-GEMM operators in performance optimization. Also, the dominant operators are diverse in each model, which indicates an optimization technique tailored for a single operator cannot fully address the non-GEMM performance problem.
Figure 5: Breaking down the execution time of models running on the Data Center (CPU only) and (CPU + GPU) configurations across operator groups.
@@ IV-B The Impact of Deployment Flow on Non-GEMM Performance
Deployment flows such as ONNX runtime [17] and TensorRT [43] are widely used for serving model inferences. Such flows apply various optimizations to each model, which includes the computational graph level optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs). To understand the effect of deployment frameworks on the operator performance, we conduct two case studies (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs(focus: general optimizations w/o operator fusion) on LLMs, and (2) comparing PyTorch and TensorRT results (focus: operator fusion).
IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
We run two LLMs (GPT2 and Llama2) using the CUDA execution provider in ORT and compare their non-GEMM performance percentage in Figure 6.
In the results, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average. We observe the percentage of memory operators significantly amplifies after applying ORT, from 3.2% to 66.8%. Such a result originates from the limitation in operator support in ORT. Many memory operators in the evaluated LLMs are not supported by the cuda execution provider in ORT, which leads to inefficient execution on CPUs involving data transfer between a CPU and a GPU. Combined with the high frequency of those operators, the inference latency of memory operators increases significantly, as shown in Figure 6 (b). The results imply two insights: (1) Model deployment flow can significantly aggravate the non-GEMM performance challenge and (2) the dominant non-GEMM operators differ depending on the operator support in a deployment flow.
@@ IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
Figure 7: The latency of non-GEMM operator in (a) Swin-b, (b) DETR, and (c) Segformer is still considerable despite applying operator fusion in TensorRT. The numbers at the top of each bar represent the latencies in milliseconds.
IV-B2 Case Study 2: Non-GEMM Performance with Operator Fusion
@@
Operator fusion is a key optimization technique for accelerating inference workloads [14, 25, 57, 42, 43, 30]. Fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42]. TensorRT [43] is a widely adopted inference framework released by Nvidia that applies the operator fusion technique targeting GPUs. Operator fusion in TensorRT detects specific patterns (e.g., three consecutive element-wise operators [43]) in the operator graph and fuses nodes captured in the patterns to enhance inference performance by reducing redundant memory accesses around non-GEMM operators.
@@
Impact of Operator Fusion. To understand the impact of operator fusion on non-GEMM performance horizon, we conduct a case study on four models, which compares the TensorRT (with fusion) and PyTorch (without fusion). We present the results in Figure 7, which shows the inference latency breakdown between GEMM and non-GEMM operators. We notice that fusion does not completely address the non-GEMM bottleneck in Swin-b model. For example, the percentage of non-GEMM operators in the total latency changes from 56.2% to 32.2%, on average, after operator fusion by TensorRT. The reduction in the percentage is based on the improvement of non-GEMM performance via operator fusion, which reduces 88.6% of latency after fusion, as summarized in Table V. However, the non-GEMM operators still account for 32.2% of total latency, which is considerable. This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance.
@@
Factors Affecting the Effectiveness of Fusion. Although most results indicate considerable impact of non-GEMM even after operator fusion, we observe operator fusion on DETR model is exceptionally effective. Therefore, we conduct a deep-dive study, investigating the percentage of fused non-GEMM operators (i.e., fusion rate) and performance improvements after fusion, as listed in Table V. We observe the strong non-GEMM performance improvements for DETR originates from high fusion rate of 30%, which led to 13.5$\times$ non-GEMM speedup. This leads to the large percentage reduction of non-GEMM in the total latency, from 67% to 18.5%, on average.
@@
However, the fusion rate is not the only factor that determines the non-GEMM speedup. For example, DETR and Segformer have similar fusion rates (30% andn 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively. We analyze the execution trace and find that batch normalization is the main source of the differences. The main difference is the fusion pattern: most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total). The results indicate that the effectiveness of operator fusion relies on not only the overall fusion rate but also the fusion patterns. Our observation confirms that the operator fusion cannot fully address the non-GEMM performance challenge, even if it can be very effective on some patterns.
IV-C Non-GEMM Performance with Quantization
Quantization refers to the model optimization technique, which reduces the bit precision of model weights and/or activations to enhance the computational performance and efficiency of DNN inference. Quantization is a widely-adopted technique [28, 40] including heavy models like LLMs [19, 35, 63, 16]. For example, LLM.int8() [16] introduced an 8-bit - 16-bit mixed precision quantization scheme for LLMs. LLM.int8() is a state-of-the-art quantization method, which quantizes more than 99% of the linear layers in OPT LLM to an 8-bit precision. Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance to understand the impact of quantization on the non-GEMM operator performance problem.
We conduct a case study on quantization using LLM.int8() on Llama3 LLM on the datacenter platform with GPU acceleration and present the results in Figure 8. In the results, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average. Such a significant shift in the latency distribution is mainly based on the GEMM performance improvements from 8-bit arithmetic, which reduced the latency by 38.2%, on average. However, based on our analysis on the execution traces, non-GEMM performance aggravates because the 8-bit data need to be dequantized and re-quantized for non-GEMM operators, which requires 16-bit floating point arithmetic. This led to 6510 additional non-GEMM operators into the computational graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization. Combined together, the overall percentage of non-GEMM operators in the total latency dominate after quantization, which makes non-GEMM operators as the major optimization target.
Figure 8: Non-GEMM operators latency contribution to the inference of a quantized Llama3 8B using 8 bit quantization [16]. The numbers at the top of each bar represent the latencies in milliseconds.
@@ IV-C Non-GEMM Performance with Quantization
TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage w.r.t the total inference latency.
@@
Swin-t 8.75% 7.53 ms (56.4%) 0.97 ms (39%)
Swin-b 7% 14.59 ms (58%) 1.65 ms (32.25%)
DETR 30% 32.17 ms (66.4%) 2.38 ms (18.5%)
Segformer 27% 5.57 ms (41%) 2.33 ms (41%)

### v5

@@ II-A ML Operators
II-B ML Models and Popular Tasks
The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of modalities and tasks (e.g. computer vision and NLP). As highlighted in  Figure 3, computer vision (Figure 3 (a) and (b)) and NLP (Figure 3 (c)) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
For example, traditional image classification models are often based on the convolutional neural network (CNN) architecture, which cascades GEMM (Conv2d) and non-GEMM (normalization and activation)operators [51, 25]. Object detection models, such as Mask R-CNN [24], often utilize CNNs for feature extraction, region proposal, and classification, as illustrated in Figure 3 (b). Unlike image classification models, they combine the CNNs with unique non-GEMM operators such as non-maximum suppression (NMS) and ROI Align to process and filter the bounding boxes for objects. On the other hand, recent language models employ the transformer architecture, which leverages the attention mechanism introduced in [57]. Transformers combine a unique set of GEMM (BMMs and Linear) and non-GEMM (normalization, memory, and element-wise arithmetic) operators, as shown in Figure 3 (c).
As we can find in the aforementioned examples, the model architectures and the combination of GEMM and non-GEMM operators are diverse. This would mean that the performance implication of non-GEMM operators would vary across models, as our motivational data presented in  Figure 1 show. This motivates a thorough characterization study that investigates (1) if the non-GEMM performance challenge is pervasive across popular models and (2) how significant their implication is, under widely adopted optimization techniques (e.g., integer quantization [29] and operator fusion [42]). Therefore, we conduct a thorough case study of the non-GEMM performance horizon.
Figure 4: An overview of NonGEMM Bench flow.
III Performance Characterization Methodology
To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets. However, ML Benchmarks available today (e.g., MLPerf [47]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators. Long-tail bench [35] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domains. Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench. NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV. To capture the performance in the latest ML workload, we select 17 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis. We discuss the models and datasets adopted in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
III-A Models included in NonGEMM Bench
Table II lists the NonGEMM Bench model registry which contains 17 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 3.7M to 7B, demonstrating the diverse model coverage of NonGEMM Bench. The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
Image Classification (IC). Image classification refers to a CV task that identifies a class label of a given image. Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features. NonGEMM Bench includes six most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [33], and three variants of Swin Transformer  [38].
Image Segmentation (IS). Image segmentation refers to a computer vision task that identifies the area in a image for each class. Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object. NonGEMM Bench includes two state-of-the-art IS models: Segformer  [61] and MaskFormer  [13].
Object Detection (OD). Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object. OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects. Using Region of Interest (RoI) processing non-GEMM operators in Table I, OD models refine raw region proposals generated by a region proposal network. The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI. NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [50], MaskRCNN  [24], and DETR  [12].
Natural Language Processing (NLP). NLP refers to tasks involving the analysis and understanding of human (natural) language. NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [64] [11]. Transformer [57] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [46] and Llama  [56].  Figure 3 (c) shows the layer architecture of GPT’s transformer. It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Computation and Element-wise Arithmetic non-GEMM operators ( Table I). NonGEMM Bench includes six popular NLP models [3]: Bert [19], three variants of GPT2 [7], Llama2-7b [56], and Mixtral 8x7B [30].
III-B NonGEMM Bench Inputs
As shown in Figure 4, NonGEMM Bench flow receives workload and dataset information (default: 17 NonGEMM Bench models in Table II), target deployment flow, and other configurations such as the batch size and number of runs for performance characterization.
Models. As described in  Section III-A, NonGEMM Bench includes a registry of 17 selected popular ML models. Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators. Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
Deployment Flow. NonGEMM Bench supports four popular inference frameworks: ONNX Runtime [49], PyTorch [44], TensorRT [4], and TorchInductor [8].
Datasets. To evaluate the models, NonGEMM Bench utilizes real datasets popular in each domain. We use ImageNet 2012  [17] and MS COCO  [37] for computer vision tasks. As for language models, we use wikitext dataset [39] available on HuggingFace. For custom models, NonGEMM Bench allows users to specify their own dataset for their models.
Configurations. Users can specify detailed configurations for the performance characterization using NonGEMM Bench. The configurations include the batch size, the number of profiling iterations, and the target device.
III-C NonGEMM Bench Outputs
NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports.
Performance Report. The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs (Figure 6) and the end-to-end energy consumption (Figure 5).
Workload Report. The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data.
Non-GEMM Report. The non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., DETR [12] employs two variant of BatchNorm, a custom implementation and one in the PyTorch operator library) and non-GEMM operator trace on different domains.
TABLE II: Tasks and Models Evaluated in Section IV
Application Models # Params Dataset
Image Classification (IC) ViT base (Vt-b) [33] 307M ImageNet  [17]
ViT large (Vt-l) [33] 307M
ViT huge (Vt-h) [33] 632M
Swin tiny (Sw-t) [38] 29M
Swin small (Sw-s) [38] 50M
Swin base (Sw-b) [38] 88M
Object Detection (OD) FasterRCNN (FRCNN) [50] 42M COCO  [37]
MaskRCNN (MRCNN) [24] 44M
DETR [12] 41M
Image Segmentation (IS) Maskformer (MF) [13] 102M COCO  [37]
SegFormer (Seg) [61] 3.7M
Natural Language Processing (NLP) GPT2 [7] 117M wikitext  [39]
GPT2 Large (gpt2-l) [7] 762M
GPT2 X-Large (gpt2-xl) [7] 1.5B
Llama 2-7B [56] 7B
Bert [19] 110M
Mixtral 8x7B [30] 46.7B
III-D NonGEMM Bench Performance Characterization Flow
NonGEMM Bench’s software flow accepts inputs described in Section III-B and generates outputs described in Section III-C. Internally, NonGEMM Bench includes graph extractor, data preprocessing, and performance profiling modules.
Graph Extractor. The Graph Extractor module extracts the operator graphs of input models based on the selected deployment flow. NonGEMM Bench utilizes graph exporters in the HuggingFace Transformers [59] and PyTorch.
Data Preprocessing. The Data Preprocessing module includes model-specific preprocessing functions that fetch raw data from the target dataset, clean the data, and apply desired transformations (e.g., tokenization and image to tensor).
Performance Profiling. The Performance Profiling (PP) module launches the inferences, collects performance statistics, and generates output reports discussed in Section III-C. The module selects appropriate profiling functions based on the deployment flow choice. For PyTorch, the PP module utilizes the PyTorch Profiler [5]. For TensorRT [6], the PP module leverages its profiling APIs. For ONNX Runtime [49], the PP module invokes Execution Provider profiling.
@@ IV Case Studies
We conduct a thorough performance analysis of models in Table II on a data center and a workstation-class platform, as listed in  Table III. We employ PyTorch for our the main performance characterization and use ONNX Runtime, TensorRT, and TorchInductor for deep-dive studies (e.g., the impact of deployment flow choice and operator fusion).
We first focuse on the GEMM and non-GEMM performance horizon (Section IV-A). Also, we provide deeper insights into the non-GEMM performance horizon by investigating the impact of deployment flow and operator fusion (Section IV-B), and the impact of quantization (Section IV-C).
@@ IV-A Non-GEMM Performance Characterization Results
Figure 5: End-to-End inference GPU energy consumption of models running on the Data Center (CPU + GPU) configuration.
We conduct a performance characterization study using PyTorch and present the results in Figure 6. Aligned with what we observed in  Figure 1, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GEMM acceleration using a GPU, from 17.2% to 42.3%, on average. However, we observe each model show different trends, mainly affected by the non-GEMM operator types and the number of GEMM and non-GEMM operators in the original model. We summarize the most time-consuming non-GEMM operators in Table IV from the data center class platform (Platform A), which shows the diversity of the dominating non-GEMM operators in each model. We highlight some notable models in each task category and delve into the details of their non-GEMM performance.
@@
IS Task: Maskformer. MaskFormer utilizes Swin Transformer as its backbone, which introduces many memory operators as we discussed for Swin Transformer. As a result, memory operator becomes the most dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 6 (f).
NLP Task: GPT2. As we observe in Figure 6 (h) for both platforms, the latency of non-GEMM operators in GPT2 variants is considerable, which account for 45.0%, on average. The most dominant non-GEMM operator is an activation function, GELU, which accounts for 26.4% of the total latency.
Summary. We observe GPU acceleration significantly increases the percentage of non-GEMM operators in the end-to-end latency, which amplifies the importance of non-GEMM operators in the performance optimization process. Also, we observe the most dominant non-GEMM operators are diverse depending on the model. The results indicate that an optimization technique tailored for a single operator cannot fully address the non-GEMM performance challenge, which motivates a holistic optimization approach for wide non-GEMM operators or a balanced specialization for a set of non-GEMM operators in a target workload.
@@
Mixtral Memory 43.1
Figure 6: Latency break-downs of NonGEMM Bench models into the operator granularity. We show CPU-only (left column) and CPU+GPU (right column) results on two evaluated platforms listed in Table III.
@@ IV-B The Impact of Deployment Flow on Non-GEMM Performance
Deployment flows such as ONNX Runtime [49] and TensorRT [4] are widely used for serving model inferences. Such flows apply various optimizations to each model, which includes the computational graph optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs). To understand the impact of deployment frameworks on the non-GEMM operator performance, we conduct two case studies: (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs (focus: general optimizations w/o operator fusion) and (2) comparing PyTorch, TorchInductor, and TensorRT results (focus: operator fusion).
[Case Study 1] Non-GEMM Performance on LLMs across PyTorch and ORT. We profile the GEMM and non-GEMM performance of two LLMs (GPT2 and Llama2) on the platform A, using the CUDA execution provider in ORT. As the results presented in Figure 7 show, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average. We observe the percentage of memory operators significantly increases if we switch from PyTorch to ORT, from 3.2% to 66.8%, although the absolute end-to-end latency decreases. Such a result originates from ORT’s significant performance boost of other operators (Lllma2) and ORT’s limited efficiency on memory operators (GPT2-XL). Many memory operators in the evaluated LLMs are not supported by the CUDA execution provider in ORT, which leads to inefficient execution on CPUs involving costly data transfer between a CPU and a GPU. Combined with the high frequency of such operators, the relative contribution of memory operators to the end-to-end inference increases significantly, as shown in Figure 7 (b). The results imply two insights: (1) Model deployment flow can significantly aggravate the non-GEMM performance challenge and (2) the dominant non-GEMM operators differ depending on the operator support in a deployment flow.
@@
Figure 8: The latency of non-GEMM operator in (a) Swin-t, (b) Swin-b, (c) DETR, and (d) Segformer is still considerable despite applying operator fusion in TorchInductor (middle) and TensorRT (right).
@@
[Case Study 2] Non-GEMM Performance with Operator Fusion. Operator fusion is one of the key optimization technique for accelerating inference workloads [16, 27, 54, 42, 4, 32]. Operator fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42]. TensorRT [4] is a widely adopted inference framework released by Nvidia that applies the operator fusion technique targeting GPUs. Operator fusion in TensorRT detects specific patterns (e.g., three consecutive element-wise operators [4]) in the operator graph and fuses nodes captured in the patterns to enhance inference performance by reducing redundant memory accesses around non-GEMM operators.
@@
To understand the impact of operator fusion on the non-GEMM performance, we conduct a case study on four models listed in Figure 8, comparing TensorRT (with fusion) and PyTorch (without fusion). We present the results in Figure 8, which shows the inference latency breakdown between GEMM and non-GEMM operators on Platform A (data center class). The results indicate that fusion mitigates the non-GEMM bottleneck, but it does not completely address the challenge. For example on Swin-b, after applying operator fusion by switching to TensorRT from PyTorch, the contribution of non-GEMM operators to the total latency changes from 56.4% to 32.2%, on average. The reduction in the percentage is based on the non-GEMM performance improvement via operator fusion, which reduces 88.6% of latency, as summarized in Table V. However, the non-GEMM operators still account for 32.2% of total latency. This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance optimization.
@@
Although most results indicate a considerable impact of non-GEMM even after operator fusion, we observe that TensorRT operator fusion on the DETR model is exceptionally effective. Therefore, we conduct a deep-dive study, investigating the percentage of fused non-GEMM operators (i.e., fusion rate) and performance improvements after fusion, as listed in Table V. We observe the strong non-GEMM performance improvements for DETR originates from high fusion rate of 30%, which led to 13.5$\times$ non-GEMM speedup. This leads to the large percentage reduction of non-GEMM in the total latency, from 66.5% to 18.5%, on average.
@@
However, the fusion rate is not the only factor that determines the non-GEMM speedup. For example, DETR and Segformer have similar fusion rates (30% and 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively. We analyze the execution trace and identify the fusion pattern around batch normalizations as the main source of the difference. Most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total). The results indicate that the effectiveness of operator fusion relies on not only the overall fusion rate but also the fusion patterns. Our observation confirms that the operator fusion cannot fully address the non-GEMM performance challenge, even if it can be very effective on some patterns.
@@
TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage with resect to the total inference latency.
@@
Swin-t 8.8% 7.53 ms (56.4%) 0.97 ms (39.0%)
Swin-b 7.0% 14.59 ms (56.4%) 1.65 ms (32.3%)
DETR 30.0% 32.17 ms (66.4%) 2.38 ms (18.5%)
Segformer 27.0% 5.57 ms (41.0%) 2.33 ms (41.0%)
IV-C The Impact of Quantization Non-GEMM Performance
Quantization refers to the model optimization technique, which reduces the bit precision of model weights and/or activations to enhance the computational performance and efficiency of DNN inference. Quantization is a widely-adopted technique [29, 40] including heavy models like LLMs [21, 36, 60, 18]. LLM.int8() is a state-of-the-art quantization method, which quantizes more than 99% of the linear layers in OPT LLM to an 8-bit precision. Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance of Llama3 on Platform A to understand the impact of quantization on the non-GEMM operator performance problem.
As the results in Figure 9 show, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average. Such a significant shift in the latency distribution is mainly based on the GEMM performance improvements from 8-bit arithmetic, which reduced the latency by 38.2%, on average. However, based on our analysis on the execution traces, non-GEMM performance aggravates because the 8-bit data need to be dequantized and re-quantized for non-GEMM operators, which requires 16-bit floating point arithmetic. This introduced 6510 additional non-GEMM operators into the computation graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization. Combined together, the overall percentage of non-GEMM operators in the total latency dominate after quantization, which makes non-GEMM operators as the major optimization target.
Figure 9: GEMM and Non-GEMM latency breakdown of the inference latency on Platform A, running an 8-bit quantized Llama3 8B [18].